3. Giovanni Briganti, the owner of "Picante Tomatoes" sells 600 kg of tomatoes at a price of $2 per kilogram. He finds that he can increase his revenue to $1296 when he sells another 48 kg. • Calculate total revenue, average revenue, and marginal revenue for this product at each output level. • Is the business an example of a perfect competitor? Explain • What is the relationship among the business's price, marginal revenue, and average revenue. What factors are responsible for this relationship? 4. What are three important decisions that a perfect competitor faces? Explain. 5. Can the government regulate monopolies? Explain. 
Answer #1
Q( units)Price( AR)TR( P*AR)MR( TRn-TRn-1)
600212002
648212962

Ans(2)---

Yes, the business is an example of perfect competition.

Reason----

Under perfect Competition,all the firms in the industry are PRICE TAKERS,that is they charge the price as determined by the industry ,as per market forces of demand and supply.

Ans(3)-----

Relationship between AR,P and MR----

Under perfect Competition,price (P) or average revenue( AR) and marginal revenue ( MR) remains same as each firm charges the same price( AR)

P or AR----- It is the market price which is determined by industry.

P= price per unit of Quantity

We find that at Q=600 or 648, the price remains same,that is $2

MR---- Marginal revenue is the addition to total revenue on selling an additional unit of output.

MR= TRn-TRn-1. Or. Change in TR/ change in Q

MR= (1296-1200)/(648-600)=$2

So, AR=MR at each level under perfect Competition.

Ans(4)----

3 IMPORTANT DECISIONS FACED UNDER PERFECT COMPETITION-----

1) Goods produced by each firm must be homogeneous .

2) All Competitive firms are price takers

3) There is no restriction in the entry and exit of firms.

Ans(5)-----

Yes ,the govt can regulate the monopolies.

It is done when the goods are essential commodities and there is a need to protect the consumer from the exploitation of producer.

The regulation is in the form of price control ,that is PRICE CEILING,under which ,the govt imposes limit on Maximum price to be charged by a monopolist.


answered by: Bhargav Agravat